### v1

# High Fidelity Text-Guided Music Generation and Editing via Single-Stage Flow Matching

Abstract
We introduce a simple and efficient text-controllable high-fidelity music generation and editing model. It operates on sequences of continuous latent representations from a low frame rate 48 kHz stereo variational auto encoder codec that eliminates the information loss drawback of discrete representations. Based on a diffusion transformer architecture trained on a flow-matching objective the model can generate and edit diverse high quality stereo samples of variable duration, with simple text descriptions. We also explore a new regularized latent inversion method for zero-shot test-time text-guided editing and demonstrate its superior performance over naive denoising diffusion implicit model (DDIM) inversion for variety of music editing prompts. Evaluations are conducted on both objective and subjective metrics and demonstrate that the proposed model is not only competitive to the evaluated baselines on a standard text-to-music benchmark - quality and efficiency-wise - but also outperforms previous state of the art for music editing when combined with our proposed latent inversion. Samples are available at https://melodyflow.github.io.
1 Introduction
Text-conditioned music generation has made tremendous progress in the past two years (Schneider et al., 2023; Huang et al., 2023; Agostinelli et al., 2023; Copet et al., 2024; Ziv et al., 2023; Liu et al., 2023; Li et al., 2023). The dominant approach involves representing audio as a sequence of compressed discrete or continuous representations and training a generative model on top of it. Two dominant generative model architectures have emerged, one based on Language Models (LMs) (Agostinelli et al., 2023; Copet et al., 2024), the other on diffusion (Schneider et al., 2023; Huang et al., 2023; Liu et al., 2023; Li et al., 2023). A third method sometimes referred to as discrete diffusion relies on non-autoregressive masked token prediction (Ziv et al., 2023; Garcia et al., 2023). The target level of audio fidelity depends on the models and some have already successfully generated 44.1 kHz or high stereo signals (Schneider et al., 2023; Li et al., 2023; Evans et al., 2024a). However no such work has really discussed the implications of aiming for higher fidelity or stereo and what performance compromise it implies from the latent representation and generation model perspective.
Due to the growing popularity of diffusion models in computer vision, a new area of research has emerged focusing on text-controlled audio editing (Wang et al., 2023; Lin et al., 2024; Garcia et al., 2023; Wu et al., 2023; Novack et al., 2024; Zhang et al., 2024; Manor and Michaeli, 2024). The creative process for sound design typically involves multiple iterations and using efficient editing methods is an effective approach to achieving this. Music editing constitutes an open ended list of tasks such as inpainting/outpainting, looping, instrument or genre swapping, vocals removal, lyrics editing, tempo control or recording conditions modification (e.g. from studio quality to a concert setting). Some of those tasks have been addressed in recent works with specialised models (Wang et al., 2023; Garcia et al., 2023; Lin et al., 2024; Wu et al., 2023; Copet et al., 2024) or zero-shot editing methods from the computer vision domain that are exclusive to diffusion models (Novack et al., 2024; Zhang et al., 2024; Manor and Michaeli, 2024). However, none of these have demonstrated high fidelity generic style transfer capabilities, due to either lack of high quality data or foundational music generation model, or design choices that do not generalise to any kind of editing task. The question of inference speed is key for creatives and the music domain is particularly challenging due to the high fidelity (48 kHz stereo) requirement in the sound design process. Recently the Flow Matching (FM) generative modeling formulation has been introduced (Lipman et al., 2022), that consists in building optimal transport paths between data and noise samples. It is a more robust and stable alternative for training diffusion models with notably faster inference and was successfully applied to train foundational speech (Le et al., 2024) and audio (Vyas et al., 2023) generative models.
In this work we present MelodyFlow, a single-stage text-conditioned FM model for instrumental music generation and editing. The model operates on continuous representations of a low frame rate Variational Audio Encoder (VAE) codec, which is crucial for handling high-fidelity stereo generation efficiently compared to models operating on discrete neural codecs. Our evaluations show that the proposed model is competitive with evaluated baselines on text-to-music generation while being remarkably efficient, with a latency of less than 10 seconds for a 30 seconds sample generation (64 inference steps). Additionally, thanks to the versatility of FM, MelodyFlow is compatible with any zero-shot test-time editing method such as DDIM inversion (Song et al., 2020). We enhance the editability of the FM inversion process with a regularized latent inversion method inspired by that of (Garibi et al., 2024) for diffusion-based image editing. Both our objective and subjective evaluations on music editing indicate that MelodyFlow can support a diversity of editing tasks on real songs without any finetuning, achieving fast music editing with remarkable consistency, text-adherence and minimal quality degradation compared with original samples. In addition we conduct an ablation study on the importance of the key design choices on the overall model quality/efficiency trade off.
Our contributions: (i) We introduce an efficient single-stage text-to-music FM model to generate 48 kHz stereo samples of up to 30 seconds, with enhancements in both the audio latent representation and generative model, striking a better balance between quality and efficiency. (ii) We present a novel regularized FM inversion method capable of performing faithful zero-shot test-time text-guided editing on various axes while maintaining coherence with the original sample. (iii) We conduct comprehensive objective and human evaluations on the key design choices underlying our approach.
2 Method
MelodyFlow is a non-causal transformer-based flow matching model conditioned on T5 text representations. It operates on sequences of latent audio representations from a neural quantizer-free audio codec with a VAE bottleneck. Combined with latent inversion, the model is able to perform text-guided editing of real or generated audio samples.
2.1 Latent audio representation
Our audio codec derives from EnCodec (Défossez et al., 2022) with additional features from the Descript Audio Codec (DAC) (Kumar et al., 2024) (snake activations, band-wise STFT discriminators) and (Evans et al., 2024a) (VAE bottleneck, perceptual weighting). A convolutional auto-encoder encodes the raw waveform into a sequence of latent bottleneck representations, its frame rate function of the convolution strides. Audio fidelity is enforced by multi-scale STFT reconstruction losses (both genuine and adversarial). The VAE bottleneck enforces smoothness in the latent representation space, as opposed to discretization bottlenecks such as Residual Vector Quantization (RVQ) (Défossez et al., 2022). Although (Vyas et al., 2023) have successfully trained a flow matching model on continuous latent representations extracted from the encoder of a discrete audio codec (before the RVQ layer), we argue that the latent space of such a codec is not the best suited for continuous modeling. Moreover while discrete representations have shown promising results on generating 32 kHz mono music data (Copet et al., 2024), adapting such models for higher fidelity stereo generation requires either scaling the vocabulary size, frame rate or number of residual projections to a point where discrete autoregressive LM becomes prohibitively more expensive than non-autoregressive continuous FM. Indeed single-stage LMs require some degree of parallel token prediction to keep the inference time acceptable (Le Lan et al., 2024). We show in section 4.3.1 that a VAE quantizer-free codec can operate at a much lower frame rate than its discrete equivalent for a better level of fidelity, reducing the cost of generation model training and inference by essentially the same factor.
2.2 Conditional flow matching model
Given an audio sample $\mathbf{a}\in\mathbb{R}^{D\times f_{s}}$, a sequence $\mathbf{x}\in\mathbb{R}^{L\times d}$ of latent representations is extracted by the neural codec. Flow matching models optimal transport paths that map a sequence $\mathbf{\epsilon}\in\mathbb{R}^{L\times d}\sim\mathcal{N}(0,I)$ to $\mathbf{x}$ trough a linear transformation - function of the flow step $t$ following equation 1.
$\mathbf{z}_{t}=t\mathbf{x}+(1-t)\mathbf{\epsilon},t\in[0,1]$ (1)
During training, $t$ is randomly sampled and the neural network $\Theta$ is trained to estimate the derivative $d\mathbf{z}_{t}/dt$ conditioned on $t$ and a text description $c$.
$d\mathbf{z}_{t}/dt=v_{\Theta}(\mathbf{z}_{t},t,c)=\mathbf{x}-\mathbf{\epsilon}$ (2)
By design, after training, the model can be used with any ODE solver to estimate $\mathbf{x}=\mathbf{z_{1}}$ given $\mathbf{\epsilon}=\mathbf{z_{0}}$ (and vice versa), and a text description. The text-to-music inference happens as such: starting from a random noise vector $\mathbf{\epsilon}\in\mathbb{R}^{L\times d}\sim\mathcal{N}(0,I)$ and a text description $c$ of the expected audio the ODE solver is run from $t=0$ to $t=1$ to estimate the most likely sequence of latents $\mathbf{x}_{generated}$.
$\mathbf{x}_{generated}=\textbf{ODE}_{0\xrightarrow{}1}(\epsilon,c)$ (3)
[…]
2.3 Text-guided editing through latent inversion
Due to the bijective nature of the FM formulation (where given a text condition, each latent sequence is mapped to a single noise vector), the model is compatible with existing latent inversion methods such as DDIM inversion (Song et al., 2020). Given the latent representation $\mathbf{x}_{orig}$ of an existing audio with an optional accompanying caption $c\in\{\varnothing,c_{orig}\}$, the model can serve to estimate its corresponding noise (or intermediate) representation $\mathbf{z}_{t_{edit}}=\textbf{ODE}_{t_{edit}\xleftarrow{}1}(\mathbf{x}_{orig},c)$ by running the ODE solver in the backward direction until an intermediary time step $t_{edit}$. It is important to note that neither $\mathbf{x}_{orig}$ nor $c_{orig}$ have to be human-generated. $\mathbf{x}_{orig}$ could be the output of a preceding text-to-music or music editing process and $c_{orig}$ could result from a music captioning model. Given the intermediary representation $\mathbf{z}_{t_{edit}}$, the ODE forward process can be conditioned on a new text description $c_{edit}$ that materialises the editing prompt: $\mathbf{x}_{edit}=\textbf{ODE}_{t_{edit}\xrightarrow{}1}(\mathbf{z}_{t_{edit}},c_{edit})$. A good inversion process should accurately reconstruct the input when $c_{edit}=c_{orig}$, as shown in equation 4.
$\mathbf{x}_{edit}=\textbf{ODE}_{t_{edit}\xrightarrow{}1}(\textbf{ODE}_{t_{edit}\xleftarrow{}1}(\mathbf{x}_{orig},c\in\{\varnothing,c_{orig}\}),c_{orig})\approx\mathbf{x}_{orig}$ (4)
In such case when swapping $c_{orig}$ for $c_{edit}$ in the $t_{edit}\xrightarrow{}1$ forward direction, the expectation is for the generated audio to preserve some consistency with the original while being faithful to the editing prompt. However in practice it was observed by (Mokady et al., 2023) that DDIM inversion suffers from poor editability due to the classifier free guidance.
2.4 Optimizing for editing
The quality of music editing depends on the codec, FM model and the latent inversion method, while we know speed is a crucial factor in the sound design creative process. Some of our design choices are critical in striking a good balance between quality and efficiency.
2.4.1 Codec VAE bottleneck
We show in section 4.3.1 that swapping the RVQ by a VAE bottleneck is key to increase the generation model efficiency and the quality of its outputs. Namely this allows better waveform reconstruction performance for a much lower frame rate, which unlocks faster inference and scaling to high-fidelity stereo audio (shown in section 4.3.2).
2.4.2 Minibatch coupling
(Tong et al., 2023; Pooladian et al., 2023) expanded over prior work on flow matching modeling by sampling pairs $(\mathbf{x},\mathbf{\epsilon})$ from the joint distribution given the by the optimal transport plan between the data $\mathbf{X}=\{{\mathbf{x}^{(i)}}\}_{i=1}^{B}$ and noise $\mathbf{E}=\{{\mathbf{\epsilon}^{(i)}}\}_{i=1}^{B}$ samples within a batch of size $B$. Essentially this translates into running the Hungarian algorithm so as to find the permutation matrix $\mathbf{P}$ that minimizes $||\mathbf{X}-\mathbf{P}\mathbf{E}||_{2}^{2}$. They demonstrate it results in straighter optimal transport paths during inference (that are closer to the theoretical linear mapping assumption between noise and data samples) and consequently offers better quality-efficiency trade offs. We shed light on the importance of mini-batch coupling in section 4.3.3 and we underline the overall benefit of the FM model design choices on its efficiency in appendix A.2.2.
2.4.3 Regularized latent inversion
(Garibi et al., 2024) recently proposed ReNoise, an iterative renoising mechanism applied at each inversion sampling step of a diffusion model. It enables building reversible diffusion trajectories with a good reconstruction/editability balance thanks to a noise regularization applied at each inversion step. We propose to follow a similar reasoning but in the flow matching setting. The theoretical grounds of flow matching are about estimating straight trajectories between noise and data samples, which helps in achieving efficient inference. However in practice those trajectories are never completely straight and naive DDIM inversion suffers from two problems.
  1. 1.
    When running the ODE solver, any pair of successive $(\mathbf{z}_{t_{1}},\mathbf{z}_{t_{2}})$ along the inversion path usually has estimated velocities $v_{\Theta}(\mathbf{z}_{t_{1}},t_{1},c)\neq v_{\Theta}(\mathbf{z}_{t_{2}},t_{2},c)$. Building a fully reversible inversion path requires estimating $\mathbf{z^{\prime}}_{t_{2}}$ such that $v_{\Theta}(\mathbf{z}_{t_{1}},t_{1},c)\approx v_{\Theta}(\mathbf{z^{\prime}}_{t_{2}},t_{2},c)$, for example using the convergence property demonstrated by (Garibi et al., 2024).
  2. 2.
    The distribution of predicted velocities tends to shift away from that of training, partially due to the classifier free guidance. One way for overcome this is to regularize them via gradient descent. Given a manually constructed $\mathbf{\tilde{z}}_{t}=\mathbf{x}t+\epsilon(1-t)$ and a real $\mathbf{z}_{t}$ along the inversion trajectory, and corresponding predictions $v_{\Theta}(\mathbf{z}_{t},t,c)$ and $\tilde{v}_{\Theta}(\mathbf{\tilde{z}}_{t},t,c)$. We arrange model predictions in 4x4 patches and compute the average Kullback-Leibler (KL) divergence $\mathcal{L}_{patchKL}$ between corresponding patches.
Our adaptation of the ReNoise inversion algorithm for flow matching is detailed in algorithm 1 and we provide a comparative analysis with DDIM inversion in sections 4.3.4 and 4.3.5. The main difference between our method and ReNoise is that given the FM formulation we only enforce the model prediction regularization, but neither de-correlation nor noise correction steps. In appendix A.2.3 we compare our approach with ReNoise by reformulating the FM prediction into a noise prediction task.
Algorithm 1 Proposed regularized FM inversion
Sequence of audio latents $\mathbf{x}$. Number of ODE backward steps $S$. Original text description $c\in\{\varnothing,c_{orig}\}$. ReNoise iteration steps $K$ with weights $\{w_{k}\}_{k=1}^{K}$, KL regularization weight $\lambda_{KL}$.
A noisy latent $\mathbf{z}_{T_{edit}}$ such that $\textbf{ODE}_{T_{edit}\xrightarrow{}1}(\mathbf{z}_{T_{edit}},c_{orig})\approx\mathbf{x}$.
$\Delta{t}\leftarrow(1-T_{edit})/{S}$
for $t=1,1-\Delta{t},\dots,T_{edit}+\Delta{t}$ do
     $\mathbf{z}^{(0)}_{t-\Delta{t}}\leftarrow\mathbf{z}_{t}$
     for $k=1,\dots,K$ do
         $\delta\leftarrow v_{\Theta}(\mathbf{z}^{(k-1)}_{t-\Delta{t}},t-\Delta{t},c)$
         if $w_{k}>0$ then
              sample $\epsilon\sim\mathcal{N}(0,I)$
              $\mathbf{\tilde{z}}^{(k-1)}_{t-\Delta{t}}\leftarrow\mathbf{x}t+\epsilon(1-t)$
              $\tilde{\delta}\leftarrow v_{\Theta}(\mathbf{\tilde{z}}^{(k-1)}_{t-\Delta{t}},t-\Delta{t},c)$
              $\delta\leftarrow\delta-\lambda_{KL}\nabla_{\delta}\mathcal{L}_{patchKL}(\delta,\tilde{\delta})$
         end if
         $\mathbf{z}^{(k)}_{t-\Delta{t}}\leftarrow\mathbf{z}_{t}-\delta\Delta{t}$
     end for
     $\mathbf{z}_{t-\Delta{t}}\leftarrow\frac{\sum_{k=1}^{K}w_{k}\mathbf{z}^{(k)}_{t-\Delta{t}}}{\sum_{k=1}^{K}w_{k}}$
end for
return $\mathbf{z}_{T_{edit}}$
3 Experimental setup
Model
MelodyFlow is a diffusion transformer of sizes 400M (small) and 1B (medium) parameters with U-shaped skip connections Bao et al. (2023). The model is conditioned via cross attention on a T5 representation (Raffel et al., 2020) computed from the text description of the music. The flow step is injected following (Hatamizadeh et al., 2023). Minibatch coupling is computed with torch-linear-assignement [^1] https://github.com/ivan-chai/torch-linear-assignment. MelodyFlow-small (resp. MelodyFlow-medium) is trained on latent representation sequences of 32 kHz mono (resp. 48 kHz stereo) segments of 10 (resp. 30) seconds, encoded at 20 Hz frame rate (resp. 25 Hz). From the codec perspective the only difference between encoding mono or stereo waveform is the number of input (resp. output) channels for the first (resp. last) convolution of the encoder (resp. decoder): 1 for mono and 2 for stereo. More details regarding audio representation and FM model training are provided in appendix A.1.
Inference and editing
For text-to-music inference we use the midpoint ODE solver from torchdiffeq with a step size of $0.03125$. Classifier free guidance of $5.0$ is chosen after grid search. For music editing we run our regularized inversion method with the euler ODE solver until $T_{edit}=0.04$ with the same classifier free guidance of $5.0$ applied in both ODE directions. Trajectory inversion estimation is run with $S=25,K=4,w_{k}=k-1$ and $\lambda_{KL}=0.2$.
3.1 Datasets
Training
Our training dataset is made of 10K high-quality internal music tracks and the ShutterStock and Pond5 music collections with respectively 25K and 365K instrument-only music tracks, totalling into 20k hours. All datasets consist of full-length music sampled at 48 kHz stereo with metadata composed of a textual description sometimes containing the genre, BPM and key. Descriptions are curated by removing frequent patterns that are unrelated to the music (such as URLs). For 32 kHz mono models the waveform is downsampled and the stereo channels are averaged.
Evaluation
For the main text-to-music generation results we evaluate MelodyFlow and prior work on the MusicCaps dataset (Agostinelli et al., 2023). We compute objective metrics for MelodyFlow and report those from previous literature. Subjective evaluations are conducted on a subset of 198 examples from the genre-balanced set. For ablations we rely on an in-domain held out evaluation set different from that of (Copet et al., 2024), made of 8377 tracks. The same in-domain tracks are used for objective editing evaluations. Subjective evaluations of edits are run on a subset of 181 higher fidelity samples from our in-domain test set with LLM-assisted designed prompts (more details in appendix A.1.3).
Metrics
We evaluate MelodyFlow using both objective and subjective metrics, following the evaluation protocol of (Kreuk et al., 2022; Copet et al., 2024) for generation. Reported objective metrics are the Fréchet Audio Distance (FAD) (Roblek et al., 2019) with VGGish embeddings (Hershey et al., 2017), the Kullback–Leibler divergence (KLD) with PASST audio encoder (Koutini et al., 2021) and CLAP [^2] https://github.com/LAION-AI/CLAP cosine similarity (Elizalde et al., 2023). For music editing evaluations we compute the average L2 distance between the original and edited latent sequences (LPAPS (Iashin and Rahtu, 2021)), FADedit between the distribution of source and edited samples and CLAPedit between the edited audio and the editing prompt. Subjective evaluations relate to (i) overall quality (OVL), and (ii) relevance to the text input (REL), both using a Likert scale (from 1 to 5). Additionally for music editing evaluations we report (iii) editing consistency (CON) where raters are asked to score the level of consistency between the editing and original samples, considering the provided editing prompt. Raters were recruited using the Amazon Mechanical Turk platform and all samples were normalized to -14dB LUFS (Series, 2011). For stereo samples objective evaluation the signal is down mixed into mono prior to metrics computation. For subjective ratings we keep the original audio format generated by each model.
4 Results
We first evaluate the model performance on text-to-music generation before investigating text-guided music editing. For text-to-music qualitative evaluations we compare MelodyFlow to three baselines: MusicGen, AudioLDM 2 and Stable-Audio. For MusicGen, AudioLDM 2 we use the available open source implementations and for Stable-Audio we use the public API (as of Wed. May 14 2024, AudioSpark 2.0 model version). For music editing qualitative evaluations we compare MelodyFlow to MusicGen-melody (specifically trained with chroma conditioning for the task of text-conditioned music editing), AudioLDM 2 with DDPM inversion (following (Manor and Michaeli, 2024)). Although Stable-Audio API offers a music editing setting, for practical reasons we were not able to submit hundreds of samples hence exclude it from our baselines.
4.1 Text-to-music generation
Text-to-music generation performance is reported in table 1. MelodyFlow achieves comparable performance with MusicGen, both lagging slightly behind Stable-Audio in terms of human preference. We do not report objective metrics on Stable-Audio as none were reported on the full MusicCaps benchmark Evans et al. (2024a). MelodyFlow achieves remarkable efficiency with only 64 inference steps.
Table 1: Comparison to text-to-music baselines. We report the original objective metrics for AudioLDM 2 and MusicGen. For subjective evaluations we report mean and CI95.
[…]
Reference - - - 3.67$\pm$0.10 4.04$\pm$0.10 -
AudioLDM 2 3.1 $1.20$ $0.31$ 2.79$\pm$0.08 3.40$\pm$0.08 208
MusicGen-small 3.1 $1.29$ $0.31$ 1500
MusicGen-medium 3.4 $1.23$ $0.32$ 3.40$\pm$0.08 3.79$\pm$0.07 1500
Stable-Audio - - - 3.67$\pm$0.08 3.89$\pm$0.07 100
MelodyFlow-small 2.8 1.27 0.33 3.27$\pm$0.08 3.83$\pm$0.08 64
MelodyFlow-medium 3.5 1.30 0.31 3.41$\pm$0.08 3.77$\pm$0.07 64
4.2 Text-guided music editing
The table 2 presents the main music editing results. We compare MelodyFlow-medium with our inversion method to AudioLDM 2 with DDPM inversion and MusicGen-melody. MelodyFlow outperforms both baselines on the quality and text-fidelity axes. Consistency-wise MelodyFlow lags slightly behind AudioLDM 2. Indeed during our listening tests we observe that AudioLDM 2 with DDPM inversion sometimes only generates a distorted version of the original track. Averaging on the three axes MelodyFlow sets a new baseline for zero-shot music editing at test-time.
Table 2: Comparison to baselines on text-guided high fidelity music editing of samples from the in-domain test set, using LLM-assisted editing prompts.
Model Method Ovl. $\uparrow$ Rel. $\uparrow$ Con. $\uparrow$ Avg. $\uparrow$
AudioLDM 2-music DDPM inv. 2.48$\pm$0.07 2.36$\pm$0.08 2.72$\pm$0.09 2.52
MusicGen-melody Chroma cond. 2.57$\pm$0.08 2.46$\pm$0.09 2.14$\pm$0.07 2.39
MelodyFlow-medium Reg. inv. 2.72$\pm$0.08 2.72$\pm$0.07 2.61$\pm$0.10 2.68
4.3 Ablations
4.3.1 Codec bottleneck and framerate
We ablate on the bottleneck choice for a fixed frame rate of 50Hz by comparing RVQ ((Copet et al., 2024) setting), VAE and identity in Table 3. From the codec perspective our results indicate optimal reconstruction performance with the identity or VAE bottlenecks followed by RVQ. The same ranking applies for SI-SDR (Le Roux et al., 2019). Regarding FM model performance we report a decreasing in-domain FAD in the favor the VAE. Ablating on the VAE codec frame rate shows that 5 Hz is comparable to 50 Hz codec with RVQ regarding codec performance. FM model performance is closer to that of 50 Hz codec with identity bottleneck. For MelodyFlow-small we chose to use 20 Hz.
Table 3: Codec bottleneck and framerate ablation for 32 kHz mono audio. Both compression and generative model performances are reported on the in-domain test set.
[…]
$\varnothing$ 50 0.35 18.5 0.68
RVQ 50 0.55 4.4 0.55
50 0.34 18.1 0.48
VAE 20 0.44 12.9 0.47
5 0.53 3.5 0.67
4.3.2 Scaling for high-fidelity and duration versatility.
The table 4 reports the impact of moving from mono to stereo with the same MelodyFlow-medium model size. We also report the effect of applying a L-shaped attention mask during model training to support durations shorter than 30 seconds during inference. For each sequence of length L, we randomly select a segment boundary within the range $[0,L]$. Positions before the boundary can only attend to themselves in the self-attention, while positions after it attend to the entire sequence. The in-domain FAD is reported for 10s and 30s generated segments. Our results indicate that the L-mask helps supporting versatile duration with no penalty on full-length segments, unlocking faster inference for shorter segments. However moving from mono to stereo only slightly affects the generative model performance, likely due to our mono-based evaluation.
Table 4: Ablation on L-mask and stereo for MelodyFlow-large. Each variant is trained on 30s audio segments encoded with a 25 Hz frame rate codec trained on 48 kHz audio.
[…]
2 0.40 12.48 0.59 0.65
1.48 0.65
1 0.39 13.34 0.49 0.59
4.3.3 Flow matching design choices
Table 5 presents the impact of the key design choices in the flow matching model training. Considering our baseline implementation was initially designed for text-to-speech (Le et al., 2024), it included an infilling objective that was meant to help the model handle variable length sequences that are inherent to the speech domain. We report the in-domain test FAD and the last validation MSEloss computed from the EMA checkpoint. No loss value is report for the baseline as the infilling objective facilitates the task. For the second line of the table the reported loss is weighted by the logit-normal probability density function at each sampled flow step to be comparable with logit-normal sampling. With all methods combined the in-domain FAD is reduced to 0.39 from 0.53 and consistent with the loss decrease, which validates our design choices.
Table 5: Flow matching model design ablation. FAD (resp. MSE) is reported on the in-domain test (resp. validation) set. Baseline is adapted from (Le et al., 2024) but retrained on our music latents.
[…]
baseline 16 24 uniform .53 -
-- infilling 16 24 uniform .50 .8596
$+$ sampling 16 24 logit-normal .44 .8484
$+$ batch coupling 16 24 logit-normal .42 .8322
$+$ wider model 18 18 logit-normal .39 .8310
(a) Editing FADedit vs. $T_{edit}$.
(b) Editing CLAPedit vs. $T_{edit}$.
(c) Editing LPAPS vs. $T_{edit}$.
Figure 1: Music editing quality as a function of the target inversion step $T_{edit}$. We report FADedit (Figure 1(a)), CLAPedit (Figure 1(b)) and LPAPS (Figure 1(c)) objective metrics.
4.3.4 Flow matching inversion
In the Figures 1(a), 1(b) and 1(c) we report music editing objective metrics as a function of $T_{edit}$, comparing naive DDIM inversion with MelodyFlow. For MelodyFlow inversion we ablate on the original text description used in the backward ODE process ($c_{T_{edit}\longleftarrow{}1}\in\{\varnothing,c_{orig}\}$). DDIM with text conditioning is not shown due to its instability.We observe that our proposed regularized inversion outperforms DDIM inversion, irrespective of the use of the text conditioning in the inversion process. Using $c_{orig}$ to condition the inversion shows better text adherence (higher CLAPedit) at the detriment of quality (higher FADedit). Consistency with the original song is generally much higher (lower LPAPS) than when using DDIM inversion, which correlates with our listening tests. The S-shaped FAD curves of the Figure 1(a) indicate an inversion target optimum around $T_{edit}=0.06$.
(a) FADedit vs. NFE.
(b) CLAPedit vs. NFE.
(c) LPAPS vs. NFE.
Figure 2: Efficiency-quality trade offs of MelodyFlow in the text-guided music editing setting, measured using objective metrics . Objective metrics (FADedit in the Figure 2(a), CLAPedit in the Figure 2(b) and LPAPS in the Figure 2(c)) indicate a sweet spot around 128 NFE.
4.3.5 Music editing efficiency
We compare DDIM with MelodyFlow inversion using $T_{edit}=0$ and several efficiency budgets. FADedit (Figure 2(a)), CLAPedit (Figure 2(b)) and LPAPS (Figure 2(c)) are plotted as a function of NFE. A good trade off between speed, quality, text-adherence and consistency is achieved at 128 NFEs.
5 Related work
Audio representation
Recent advancements in neural codecs have seen the application of VQ-VAE on raw waveforms, incorporating a RVQ bottleneck as demonstrated in Zeghidour et al. (2021); Défossez et al. (2022), later refined as per Kumar et al. (2024). (Evans et al., 2024a) proposed a modification to this approach by replacing the RVQ with a VAE bottleneck to enhance the modeling of continuous representations. In addition, several recent audio generative models have adopted Mel-Spectrogram latent representations, coupled with a vocoder for reconstruction, as shown in the works of (Ghosal et al., 2023; Liu et al., 2023; Le et al., 2024). Furthermore, Défossez et al. (2022) introduced an additional layer of complexity by incorporating quantization to support discrete representation on top of the continuous representation.
Text-to-music generation
Models that operate on discrete representation are presented in the works of (Agostinelli et al., 2023; Copet et al., 2024; Ziv et al., 2023). Agostinelli et al. (2023) proposed a representation of music using multiple streams of tokens, which are modeled by a cascade of transformer decoders conditioned on a joint textual-music representation (Huang et al., 2022b). Copet et al. (2024) introduced a single-stage language model that operates on streams of discrete audio representations, supporting both 32 kHz mono and stereo. Ziv et al. (2023) replaced the language model with a masked generative single-stage non-autoregressive transformer. Schneider et al. (2023); Huang et al. (2023); Liu et al. (2023) use diffusion models. Schneider et al. (2023) utilized diffusion for both the generation model and the audio representation auto-encoder. Liu et al. (2023) trained a foundational audio generation model that supports music with latent diffusion, conditioned on autoregressively generated AudioMAE features (Huang et al., 2022a). Evans et al. (2024a, b) proposed an efficient long-form stereo audio generation model based on the latent diffusion of VAE latent representations. This model introduced timing embeddings conditioning as a method to better control the content and length of the generated music.
Music editing
Lin et al. (2024) proposed a parameter-efficient fine-tuning method for autoregressive language models to support music inpainting tasks. Garcia et al. (2023) developed a masked acoustic modeling approach for music inpainting, outpainting, continuation and vamping. Wu et al. (2023) fine-tuned a diffusion-based music generation model with melody, dynamics and rhythm conditioning. Novack et al. (2024) is a fine-tuning free framework for controlling pre-trained text-to-music diffusion models at inference-time via initial noise latent optimization. Zhang et al. (2024) investigated zero-shot text-guided music editing with conditional latent space and cross attention maps manipulation. Manor and Michaeli (2024) employs DDPM inversion (Huberman-Spiegelglas et al., 2023) for zero-shot unsupervised and text-guided audio editing.
6 Discussion
Limitations
Under our current setup text-guided music editing prompts are not instructions. They describe what the edited sample should sound like given an original music sample and description, but the model is not designed to understand direct editing instructions like replace instrument A by instrument B. While we observed that MelodyFlow performs convincing editing tasks for several axes (genre or instrument swap, tempo modification, key transposition, inpainting/outpainting), more research work is required to accurately evaluate each of those axes. Music editing human listening tests are conducted for a fixed $T_{edit}$, but eventually it should depend on the sound designer’s preference on the creativity-consistency axis. Finally the reported objective metrics are mostly used as a proxy for subjective evaluations but they have their limitations. As an example we observe that optimizing FAD for MusicCaps is usually achieved by overfitting on our training dataset, which negatively correlates with perceived quality.
Broader Impact
Such a model has the potential to help creators expand their creativity by generating different styles of high quality music. The combination of efficiency and quality may eventually revolutionise the music sound design and composition work. This should unlock the ability for any user to listen to the music they like under a different style. However such a model can also give malicious users the ability to alter and republish any copyrighted content without the original creators content.
Conclusion
In this work we presented MelodyFlow, the first non-autoregressive model tailored for zero-shot test-time text-guided editing of high-fidelity stereo music. In the text-to-music setting the model offers competitive performance thanks to a low frame rate VAE codec and FM model featuring logit-normal flow step sampling, optimal-transport minibatch coupling and L-shaped attention mask. Combined with our proposed regularized latent inversion method, MelodyFlow outperforms previous zero-shot test-time methods by a large margin. The model achieves remarkable efficiency that is key for the sound design creative process and supports variable duration samples. Our extensive evaluation, that includes objective metrics and human studies, highlights MelodyFlow promise for efficient music editing with remarkable consistency, text-adherence and minimal quality degradation compared with the original, while remaining competitive on the task of text-to-music generation. For future work we intend to explore how to accurately evaluate specific editing axes and how such a model could help design metrics that better correlate with human preference.
Appendix A Appendix
A.1 Experimental setup
A.1.1 Audio latent representation
Our compression model implementation is that of Copet et al. [2024] enhanced by band-wise discriminators and snake activations from [Kumar et al., 2024], perceptual weighting [Wright and Välimäki, 2019], VAE bottleneck and multi resolution STFT reconstruction loss from [Evans et al., 2024a]. We train a mono 32 kHz codec at 20 Hz frame rate and another one supporting stereo 48 kHz audio at 25 Hz. The bottleneck dimension is of 128. Both are trained on one-second random audio crops for 200K steps, with a constant learning rate of 0.0003, AdamW optimizer and loss balancer of [Défossez et al., 2022]. Stereo codecs are trained with sum and difference loss [Steinmetz et al., 2020]. The bottleneck layer statistics are tracked during training (dimension-wise) for normalization prior to FM model training.
A.1.2 Flow matching model
MelodyFlow is a diffusion transformer that follows Esser et al. [2024] configurations where each head dimension is of 64 and the model has the same number of heads and layers (either 18 or 24). Model implementation is that of audiocraft [^3] https://github.com/facebookresearch/audiocraft but adapted for FM following [Vyas et al., 2023]: U-shaped skip connections are added along with linear projections applied after concatenation with each transformer block output Bao et al. [2023]. The model is conditioned via cross attention on a T5 representation [Raffel et al., 2020] computed from the text description of the audio, using 20% dropout rate during training in anticipation for the classifier free guidance applied at inference. Cross attention masking is used to properly adapt to the text conditioning sequence length of each sample within a batch and we use zero attention for the model to handle unconditional generation transparently. No prepossessing is applied on text data and we only rely on the descriptions (additional annotations tags such as musical key, tempo, type of instruments, etc. are discarded, although they also sometimes appear in the text description). The flow step is injected following [Hatamizadeh et al., 2023]. Minibatch coupling is computed with torch-linear-assignement [^4] https://github.com/ivan-chai/torch-linear-assignment. MelodyFlow-small (resp. MelodyFlow-medium) is trained on latent representation sequences of 32 kHz mono (resp. 48 kHz stereo) segments of 10 (resp. 30) seconds, encoded at 20 Hz frame rate (resp. 25 Hz). MelodyFlow-small (resp. MelodyFlow-medium) is trained for 240k (resp. 120k) steps with AdamW optimizer ($\beta_{1}=0.9$, $\beta_{2}=0.95$, weight decay of 0.1 and gradient clipping at 0.2), a batch size of 576 and a cosine learning rate schedule with 4000 warmup steps. Additionally, we update an exponential moving average of the model weights ever 10 steps with a decay of 0.99. Each model is trained on 8 H100 96GB GPUs with bfloat16 mixed precision and FSDP [Zhao et al., 2023]. MelodyFlow-small requires 3 days and MelodyFlow-medium 6 days of training.
A.1.3 LLM-assisted editing prompt generation
For editing prompts design we utilized the LLama-3 large language model to modify the original descriptions by targeting genre swapping. Edited descriptions were then manually verified to ensure their plausibility and coherence. As an example, given the original description This is a lush indie-folk song featuring soaring harmony interplay and haunting reverb-y harmonica, the resulting editing prompt is This is a lush Indian classical-inspired song featuring soaring harmony interplay and haunting reverb-y bansuri flute.
A.1.4 Subjective evaluation form
A screenshot of the music subjective evaluation form is shown in the Figure 3.
Figure 3: Music editing subjective evaluation form. Given the original song A, raters are asked to evaluate three different edits of A, on the three following axes: quality, text adherence, consistency.
A.2 Additional experiments
A.2.1 Classifier-free guidance
In the Figure 4(a) we report the in-domain evaluation FAD as a function of the classifier-free guidance factor in the text-to-music generation setting. Throughout the paper we use a classifier-free guidance factor of $5.0$.
A.2.2 Inference efficiency
In the Figure 4(b) we present the measured FAD as a function of inference steps for both the baseline and final version of MelodyFlow. Not only is MelodyFlow achieving better performance overall but it is able to outperform the baseline with 16 times fewer NFE.
(a) Classifier-free guidance.
(b) Generation FAD vs. NFE.
Figure 4: Text-to-music generation quality (FAD) as a function of the classifier-free guidance factor (Figure 4(a)) and inference steps (Figure 4(b)). The baseline of the Figure 4(b) is the FM model architecture of [Le et al., 2024] but retrained on our music latents. The combination of our flow matching design choices enable faster generation for a given efficiency budget or better overall quality.
A.2.3 Regularized inversion ablation
During latent inversion the FM model predicts the velocity $v_{\Theta}(\mathbf{z}_{t},t,c)=\mathbf{x}-\mathbf{\epsilon}$. The predictions are regularized using the Algorithm 1 using a weighted KL patch-wise divergence loss $\mathcal{L}_{patchKL}$. In the Figures 5(a) and 5(b) we ablate on the divergence loss weight $\lambda_{KL}$ for $T_{edit}=0.04$, using $S=25,K=4,w_{k}=k-1$. Since during latent inversion we know the original latent $\mathbf{x}_{orig}$, the model prediction can be rewritten as $\epsilon$-prediction by computing $\epsilon=v_{\Theta}(\mathbf{z}_{t},t,c)-\mathbf{x}_{orig}$. In that scenario we can apply the exact ReNoise inversion method of [Garibi et al., 2024], using 10 iterations of noise de-correlation (shown as $\epsilon$-prediction in the Figures). Whether the model prediction is expressed as noise or velocity, the Figures show that an optimum can be achieved around $\lambda_{KL}=0.15$ for velocity prediction and around $0.05$ for noise prediction. Overall the quality is better (lower FADedit in the Figure 5(a)) when directly regularizing the velocity prediction. In both cases we observe a higher CLAPedit in the Figure 5(b) when the original text description $c_{orig}$ conditions the inversion process, confirming better text-adherence. This happens at the expense of a higher FADedit compared with unconditional inversion.
(a) FADedit as a function of $\lambda_{KL}$.
(b) CLAPedit as a function of $\lambda_{KL}$.
Figure 5: Effect of the loss regularization weight $\lambda_{KL}$ on the quality (Figure 5(a)) and text-adherence (Figure 5(b)) of music editing. Noise and velocity prediction are compared, with or without the original text description $c_{orig}$.